OpenStudy (anonymous):

What is the exact volume of a cylinder with a radius of 4 inches and a height of 8 inches?

OpenStudy (igreen):

V = πr^2h

OpenStudy (igreen):

\[V = (3.14)(4)^2(8)\]

OpenStudy (igreen):

\[V = (3.14)(16)(8)\]

OpenStudy (igreen):

@QueenElsa5 Can you multiply those?

OpenStudy (anonymous):

Yeah.

OpenStudy (igreen):

Okay, what did you get?

OpenStudy (anonymous):

I got 401.92

OpenStudy (igreen):

Right. So \(\ V = 401.92\)

OpenStudy (igreen):

Nice work. You can't really get an EXACT answer because, as you know, pi goes on forever..

OpenStudy (anonymous):

Thank you.